# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ V-C5 Bounded Plan Impossibility
Consider a CT node $N\in\varsigma$ mapping to $\tau^{\infty}$. At $N$, a plan $\pi\in N_{\Pi}$ with $\pi\sim\tau^{\infty}$ is branched on. To satisfy the non-termination condition (I), $\pi_{D}\leq c$ for some fixed $c<\infty$ which requires that every timed action in $\pi$ (except the final infinite wait action, implicitly omitted in the following) ends before $c$. For simplicity, we suffice with the less restrictive requirement that all timed move actions must begin before $c$. In other words, under the constraints $N_{C}$ and $\pi_{D}\leq c$ it holds that
@@
In Section V-C3, we showed that when our branching rule is applied to a move action $\langle m,t\rangle\in\pi$, then either a non-degenerate prefix or the entirety of a maximally connected interval $s_{m},e_{m})\subseteq P_{m}$ is removed from $P_{m}$. Similarly, it was shown in Section V-C4 that when our branching rule is applied to a wait action $\langle w,t\rangle\in\pi$, then one of two cases occur. In the first case, the same constraining as when our branching rule is applied to a move action occurs for the move action following $\langle w,t\rangle$ in $\pi$. In the second case, a non-degenerate reduction of $\mathcal{T}_{v(w)}$ occurs.
Before proceeding, we analyze the one case which has yet to be addressed: when the action $\langle w_{n},t_{n}\rangle$ in $\pi\sim\tau^{\infty}$ that is branched on at $N$ is an infinite wait action, $w_{n,D}=\infty$. When our branching rule is applied to $\langle w_{n},t_{n}\rangle$, a constraint is created forbidding the use of $v(w_{n})$ during some interval $t^{\prime},t^{\prime\prime})$. From Section V-C4 we know that $t_{n}<t^{\prime\prime}$ by some non-degenerate amount. Thus, the earliest time the agent can arrive at $v(w_{n})$ to remain there indefinitely is at the later time $t^{\prime\prime}$ than $t_{n}$. Consequently, the timed move action $\langle m_{n-1},t_{n-1}\rangle\in\pi$ preceding $\langle w_{n},t_{n}\rangle$ is no longer valid since it leads to the agent arriving at $v(w_{n})$ before $t^{\prime\prime}$. Instead, $m_{n-1}$ can only occur at some time later than $t^{\prime\prime}-m_{n,D}$. Since $m_{n-1}$ is the last move action in $\tau^{\infty}$, these constraint additions lead to a non-degenerate increase in the duration of all plans $\pi^{\prime}\sim\tau^{\infty}$.
From the above, at every branching there occurs a non-degenerate reduction in one of the finite sets $P_{m}$ or $\mathcal{T}_{v}$ (with $m\in\tau^{\infty}$ and $v\in\mathbf{v}$) or a non-degenerate increase the in duration of all valid plans $\pi^{\prime}\sim\tau^{\infty}$. Therefore, after a finite many branchings either one of the sets will be depleted or the duration of valid plans mapping to $\tau^{\infty}$ will grow beyond the fixed $c<\infty$. In either case, eventually after a finite many branchings no plan $\pi\sim\tau^{\infty}$ satisfying the accumulated constraints $N_{C}$ and $\pi_{D}<c$ will exist.
To conclude, the requirement of non-termination which was established in Section V-B cannot be satisfied by CCBS under our branching rule. Therefore, termination is guaranteed after a finite number of iterations.
V-D Soundness and Solution Completeness
Theorem V.2 (Sound and solution complete).
CCBS under our branching rule is sound and solution complete.
@@ V-D Soundness and Solution Completeness
Our branching rule was shown in Section V-A to be sound. By Theorem III.2, CCBS is therefore sound when using a sound branching rule. Additionally, Section V-C proved that CCBS under our branching rule terminates within finite iterations. Since CCBS under our branching rule is both sound and guaranteed to terminate, Theorem III.3 establishes that it is solution complete. ∎
VI Experimental Evaluation
We solve the MAPFR problem in Figure 9 for minimal sum-of-costs using two variants of the publicly available CCBS implementation  [^3] https://github.com/PathPlanning/Continuous-CBS.: the unmodified CCBS release using the implemented branching rule, and an otherwise identical version using our branching rule  [^4] https://github.com/Adcombrink/S-and-SC-CCBS.. Experiments are run on a Mac Studio, 16-core M4 Max CPU, 64 GB RAM, macOS Sequoia (15.3). We use $\gamma=0.5$ when calculating $\delta$.
Figure 9: A MAPFR problem with four agents, each agent $i$ starts at vertex $S(i)$ and must reach its goal vertex $G(i)$.
The example in Figure 9 contains four agents, $i=1,..,4$, each with a start vertex $S(i)$ and goal vertex $G(i)$. All agents are circular with radius $r=\sqrt{2}/4$ and traverse edges at constant speed $1$ in a straight line. Agent $1$ starts at vertex $S(1)=E$ and must pass vertex $F$ to reach its goal $G(1)=G$, therefore, agent $2$ with $S(2)=G(2)=F$ must move out of the way. The solution found by both methods is for agent $2$ to move to $C$ to allow agent $1$ to pass. However, agent $2$ must avoid a collision with agent $3$ which passes vertex $C$ on its way to its goal vertex $D$. So, there are two options for the interaction between agents $2$ and $3$ at $C$:
@@ VI Experimental Evaluation
Due to the difference in length of edges $EF$ and $AB$, the time that agent $4$ must wait for agent $3$ is larger than the time that agent $1$ must wait for agent $2$. Therefore, these two possible solutions have different sum-of-costs, only the one where agent $2$ waits for agent $3$ is optimal.
We now examine how CCBS under the implemented branching rule handles this problem. At the CT root $N^{R}$, with $N^{R}_{C}=\varnothing$, each agent’s initial plan is generated without collision avoidance:
@@
where $\langle VU,t\rangle$ denotes a timed move action along edge $(V,U)$ starting at time $t$, and $\langle V,t_{1},t_{2}\rangle$ is a timed wait action at vertex $V$ from $t_{1}$ to $t_{2}$. Two conflicts appear in $N^{R}_{\Pi}$: $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ and $\langle\langle FG,1.5\rangle,\langle F,0,\infty\rangle\rangle$ (between agents $1$ and $2$). CCBS branches on the first conflict, with collision interval $I^{c}=[0.793,1.5)$. This yields the constraints $\langle 1,EF,[0,\infty)\rangle$ (forbidding agent $1$ from ever traversing $EF$) and $\langle 2,F,I^{c}\rangle$ (forbidding agent $2$ from occupying $F$ during $I^{c}$). The first constraint is infeasible, as agent $1$ must traverse $EF$ to reach its goal $G$. Thus, agent $2$ must vacate $F$ before $t=0.793$. However, the earliest time that agent $2$ can traverse $FC$ while allowing agent $3$ to pass first is at $t=1$, which is not possible due to the constraint. Therefore, all valid solutions where agent $2$ waits for agent $3$ are removed from the search. Consequently, agent $3$ must wait for agent $2$ to complete its detour via $C$. The resulting solution is:
@@
We now examine how CCBS under our branching rule handles the same conflict. The conflict $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ with $I^{c}=[0.793,1.5)$ has an intersection interval, i.e., the time during which agent $1$ executing $\langle EF,0\rangle$ is in collision with agent $2$ waiting at vertex $F$, of $\bar{I^{c}}=I^{c}=[0.793,1.5)$). This yields
$\delta=\min\left(0.5\,(1.5-0.793),\infty\right)=0.354$
@@
Unlike the constraints produced by standard CCBS, these restrictions exclude only infeasible behaviors while preserving feasible solutions — including the case where agent $2$ waits until after agent $3$ passes vertex $C$. The resulting solution is:
@@
The performance using each branching rule is reported in Table I. There, we see that our branching rule yields a solution with lower sum-of-costs, makespan, and computation time than standard CCBS. Consequently, this provides additional evidence to that in [28] that CCBS is not sound. Additionally, with the support of the theoretical results in Section V, the solution found by CCBS under our branching rule has optimal sum-of-costs.
@@
TABLE I: Performance comparison between standard CCBS and CCBS using our branching rule.
Branching rule Sum-of-Costs Makespan Computation Time
Standard $10.707$ $3.500$ $8.1$ ms
Ours $9.000$ $3.000$ $1.4$ ms
VII Conclusions
@@
This work set out to resolve the recently revealed long-standing gap between the theory and practice of Continuous-time Conflict Based Search (CCBS) for optimal Multi-Agent Path Finding in continuous time (MAPFR). We revisited the MAPFR formulation and analyzed the two existing branching schemes — the theoretical rule and the publicly available implementation — through a new analytical framework. This analysis yielded simple, sufficient conditions for soundness and solution completeness and showed that the implemented branching rule violates them, thereby supporting the counter-examples that cast doubt on CCBS.
We then presented a our novel branching rule and proved that it meets all conditions: it never excludes solutions from the search and guarantees termination when used in CCBS. Consequently, the resulting CCBS variant is both sound and solution complete, matching the guarantees of discrete-time CBS for the first time in the continuous domain. We validated these results by presenting an example MAPFR problem and applying standard CCBS and CCBS using our branching rule. Results shows that our branching rule produces a solution with lower sum-of-costs than standard CCBS, thereby showing definitively that standard CCBS is not sound.
Practically, our rule can be adopted as a drop-in replacement in existing CCBS code-bases to obtain certificates of optimality without risking non-termination. Conceptually, the constraint-set perspective and non-termination criterion offer tools for analyzing other CCBS-like MAPFR solvers and extensions, laying a solid foundation for the next generation of multi-agent path finding algorithms.

### v2

@@ 5.3.5 Bounded Plan Impossibility
Consider a CT node $N\in\varsigma$ mapping to $\tau^{\infty}$, meaning that a plan $\pi\in N_{\Pi}$ with $\pi\sim\tau^{\infty}$ is branched on. To satisfy the non-termination condition (I), $\pi_{D}\leq c$ for some fixed $c<\infty$ which requires that every timed action in $\pi$ (except the final infinite wait action, which we will momentarily address specifically) ends before $c$. For simplicity, we suffice with the less restrictive requirement that all timed move actions must begin before $c$. In other words, under the constraints $N_{C}$ and $\pi_{D}\leq c$ it holds that
@@
Before proceeding, we address the case of $\delta$-BR being applied to an infinite wait action $\langle w^{j}_{n},t^{j}_{n}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j}_{n},t^{j}_{n}\rangle\rangle$, with $\pi_{j}\sim\tau^{\infty}$. Let this conflict have intersection interval $\bar{I^{c}}$. $\delta$-BR creates a vertex constraint $\langle j,v(w^{j}_{n}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and several motion constraints (which are not necessary to consider for this analysis). Thus, to satisfy the vertex constraint and also follow a plan mapping to $\tau^{\infty}$ which ends with waiting at $v(w^{j}_{n})$ indefinitely, agent $j$ can only arrive at $v(w_{n})$ from time $\bar{t_{2}^{c}}$ and later. By (14) we know that $t_{n}^{j}<\bar{t_{2}^{c}}$. Consequently, the timed move action $\langle m^{j}_{n-1},t^{j}_{n-1}\rangle\in\pi_{j}$ preceding $\langle w^{j}_{n},t^{j}_{n}\rangle$ is no longer valid since it leads to the agent arriving at $v(w_{n})$ at $t^{j}_{n}$, that is, before $\bar{t_{2}^{c}}$. Instead, $m^{j}_{n-1}$ can only occur at some time later than $\bar{t_{2}^{c}}-m^{j}_{n,D}$. Since $m^{j}_{n-1}$ is the last move action in $\tau^{\infty}$, this branching leads to a non-degenerate increase in the duration of all plans $\pi^{\prime}\sim\tau^{\infty}$.
In Section 5.3.3, we showed that when $\delta$-BR is applied to a move action $\langle m,t\rangle\in\pi$, then either a non-degenerate prefix or the entirety of a maximally connected interval $s_{m},e_{m})\subseteq P_{m}$ is removed from $P_{m}$. Similarly, it was shown in Section 5.3.4 that when $\delta$-BR is applied to a wait action $\langle w,t\rangle\in\pi$, then one of two cases occur. In the first case, the same constraining as when $\delta$-BR is applied to a move action occurs for the move action following $\langle w,t\rangle$ in $\pi$. In the second case, a non-degenerate reduction of $\mathcal{T}_{v(w)}$ occurs.
From the above, at every branching there occurs a non-degenerate reduction in one of the finite safe interval sets $P_{m}$ or $\mathcal{T}_{v}$ (with $m\in\tau^{\infty}$ and $v\in\mathbf{v}$) or a non-degenerate increase in the duration of all valid plans $\pi^{\prime}\sim\tau^{\infty}$ (when a final infinite wait action is branched on). Therefore, after a finite number of iterations, either one of the safe interval sets $P_{m}$ or $\mathcal{T}_{v}$ will be depleted or the duration of valid plans mapping to $\tau^{\infty}$ will grow beyond the fixed $c<\infty$. In either case, eventually — after a finite number of branchings — no plan $\pi\sim\tau^{\infty}$ exists that can satisfy both the accumulated constraints from these branchings and the plan duration bound $\pi_{D}<c$.
To conclude, the requirement of non-termination which was established in Section 5.2 cannot be satisfied by CCBS using $\delta$-BR. Therefore, termination is guaranteed after a finite number of iterations.
5.4 Soundness and Solution Completeness
Based on the results from the previous sections, we finalize our proof that CCBS using $\delta$-BR is sound and solution complete.
Theorem 5.2 (Sound and solution complete).
CCBS using $\delta$-BR is sound and solution complete.
@@ 5.4 Soundness and Solution Completeness
$\delta$-BR was shown in Section 5.1 to be sound, which means by Theorem 3.2 that CCBS using $\delta$-BR is sound. Additionally, Section 5.3 proved that CCBS using $\delta$-BR terminates within finite iterations. With these soundness and termination guarantees, Theorem 3.3 establishes that CCBS using $\delta$-BR is solution complete. ∎
6 Experimental Evaluation
We perform experiments with the publicly available CCBS using the IBR (CCBS-IBR), and an otherwise identical CCBS using the proposed $\delta$-BR (CCBS-$\delta$-BR). Both versions are available in our repository [^3] https://github.com/Adcombrink/Optimal-Continuous-CBS. along with details in A regarding necessary modifications to the original code base. A counterexample is introduced in Section 6.1 for which CCBS-$\delta$-BR produces an optimal solution and CCBS-IBR returns a sub-optimal solution, along with explanations as to why this occurs. Thereafter, we compare the solution quality and runtime for the two methods on benchmark problems in Section 6.2. Finally, in Section 6.3 we perform an ablation study on the parameter $\gamma$ used in $\delta$-BR, motivating the use of $\gamma=0.9$ unless stated otherwise. We use circular agents with radius $\sqrt{2}/4$ that traverse edges in straight lines at constant speed $1$. All experiments are run on a 2025 Mac Studio, 16-core M4 Max CPU, 64 GB RAM, macOS Sequoia (15.3).
6.1 A Counterexample
Figure 9: A MAPFR problem with four agents with radius $r=\sqrt{2}/4$. Each agent $i=1,..,4$ is shown at its start vertex $S(i)$ and must reach its goal vertex $G(i)$ without colliding with any other agent.
The example in Figure 9 contains four agents, $i=1,..,4$, each with a start vertex $S(i)$ and goal vertex $G(i)$. All agents are circular with radius $r=\sqrt{2}/4$ and traverse edges at constant speed $1$ in a straight line. Agent $1$ starts at vertex $S(1)=E$ and must pass vertex $F$ to reach its goal $G(1)=G$, therefore, agent $2$ with $S(2)=G(2)=F$ must move out of the way. The solution found by both CCBS-IBR and CCBS-$\delta$-BR is for agent $2$ to move to $C$ to allow agent $1$ to pass. However, agent $2$ must avoid a collision with agent $3$ which passes vertex $C$ on its way to its goal vertex $D$. So, there are two options for the interaction between agents $2$ and $3$ at $C$:
@@ 6.1 A Counterexample
Due to the difference in length of edges $EF$ and $AB$, the time that agent $4$ must wait for agent $3$ is larger than the time that agent $1$ must wait for agent $2$. Therefore, these two possible solutions have different sum-of-costs, and only the one where agent $2$ waits for agent $3$ is optimal.
We now examine how CCBS-IBR handles this problem. At the CT root $N^{R}$, with $N^{R}_{C}=\varnothing$, each agent’s initial plan is generated without collision avoidance:
@@
where $\langle VU,t\rangle$ denotes a timed move action from $V$ to $U$ starting at time $t$, and $\langle V,t_{1},t_{2}\rangle$ is a timed wait action at vertex $V$ from $t_{1}$ to $t_{2}$. Two conflicts appear in $N^{R}_{\Pi}$: $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ and $\langle\langle FG,1.5\rangle,\langle F,0,\infty\rangle\rangle$, both between agents $1$ and $2$. CCBS-IBR branches on the first conflict, with intersection interval $\bar{I^{c}}=[0.793,1.5)$. This yields the constraints $\langle 1,EF,[0,\infty)\rangle$ (forbidding agent $1$ from ever traversing $EF$) and $\langle 2,F,\bar{I^{c}}\rangle$ (forbidding agent $2$ from occupying $F$ during $\bar{I^{c}}$). The first constraint is infeasible, as agent $1$ must traverse $EF$ to reach its goal $G$. Thus, agent $2$ must vacate $F$ before $t=0.793$. However, the earliest time that agent $2$ can traverse $FC$ while allowing agent $3$ to pass first is at $t=1$, which is not possible due to the constraint. Therefore, all valid solutions where agent $2$ waits for agent $3$ are removed from the search. Consequently, agent $3$ must wait for agent $2$ to complete its detour via $C$. The resulting solution is:
@@
We now examine how CCBS-$\delta$-BR handles the same conflict. The conflict $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ with intersection interval $\bar{I^{c}}=[0.793,1.5)$) yields
$\delta=\min\left(0.5\cdot\,|[0.793,1.5)|,\infty\right)=0.354$
@@
Unlike the constraints produced by CCBS-IBR, these restrictions exclude only infeasible behaviors while preserving feasible solutions — including the case where agent $2$ waits until after agent $3$ passes vertex $C$. The resulting solution is:
@@
The performance using each branching rule is reported in Table 1. There, we see that CCBS-$\delta$-BR yields a solution with lower sum-of-costs, makespan, and computation time than CCBS-IBR. Consequently, this provides additional evidence to that in [29] and Section 4 that CCBS-IBR is not sound. Moreover, the theoretical results in Section 5 ensure that the solution found by $\delta$-BR has optimal sum-of-costs. However, we further strengthen this claim by validating the optimality of CCBS-$\delta$-BR’s solution by constructing a tailor-made Satisfiability Modulo Theory (SMT) model for this specific problem and solving it using the SMT solver Z3 [33], confirming after roughly $26$ hours of computation that the solution returned by $\delta$-BR is indeed optimal. Details regarding this validation can be found in B, and animations of CCBS-IBR and CCBS-$\delta$-BR’s solutions can be found in the provided repository.
@@
Table 1: Performance comparison between the CCBS-IBR and CCBS-$\delta$-BR on the example in Figure 9.
Method Sum-of-Costs Makespan Computation Time
CCBS-$\delta$-BR $9.000$ $3.000$ $1.0$ ms
CCBS-IBR $10.707$ $3.500$ $7.1$ ms
6.2 Benchmark Comparison
We compare CCBS-IBR and CCBS-$\delta$-BR on the same $2^{k}$-neighborhood gridmaps and roadmaps benchmark sets from [21], and additionally on our own randomly generated gridlike roadmaps. The benchmarking scheme is the same as in [21]: for each map, a scenario defines a list of start and goal vertex pairs. We then create a problem with $n$ agents using the first $n$ start and goal vertex pairs. For each map and scenario, we begin by solving for $n=2$, and repeatedly incrementing $n$ until a problem cannot be solved within a time limit of $30$ seconds.
The $2^{k}$-neighborhood gridmaps — originally introduced in [8] — define a grid of traversable and blocked cells. Each traversable cell is connected to its $2^{k}$ neighborhood, provided that a traversal can be done without colliding with a blocked cell. We test with the same parameters as in [21]: k=1,,5k=1,...,5 with each of the maps Den520d, Warehouse, Rooms, and Empty16x16. The roadmaps are based on the overall shape of Den520d with three levels of density: Sparse, Dense, and Mega-dense. We refer to [21] for additional details regarding these maps.
In addition to the gridmaps and roadmaps, we introduce gridlike roadmaps which are generated with a desired average degree over the map vertices. Two example maps with average degrees of $2.0$ and $3.5$ are shown in Figure 10. By lowering the average degree, the map becomes more constrained in the number of paths between two vertices. The likelihood of agent interactions consequently increases. These maps are generated from a 2020–by–$10$ vertex grid with spacing 11 by applying zero-mean gaussian noise with standard deviation $0.15$ to the vertex positions, and randomly removing edges until a desired average vertex degree is achieved. Edges are only removed if the resulting graph remains connected.
(a) Average vertex degree $3.5$.
(b) Average vertex degree $2.0$.
Figure 10: Two gridlike roadmaps with different average degrees and agent start positions.
On the gridmap and roadmap benchmark sets, we find no difference in solution quality between CCBS-IBR and CCBS-$\delta$-BR; both methods produce equivalent solutions. However, on the gridlike roadmap benchmark set we find on $11$ problems that CCBS-$\delta$-BR produces solutions with a lower sum-of-costs than CCBS-IBR, and equivalent solutions on all remaining problems. Although these $11$ problems represent only a fraction of the 15 46215\,462 problems that both methods could solve within the time limit, it shows that CCBS-IBR’s sub-optimality issues do not only present themselves in tailor-made problems. These results also hint at the difficulty of producing benchmark problems that fully test the limits of these algorithms. Additionally, the $11$ problems that trigger CCBS-IBR to return sub-optimal solutions are on maps with an average node degree distribution toward the lower end of the tested range: $1$ with average degree $2.1$, 33 with average degree $2.2$, 44 with average degree $2.4$, and $3$ with average degree $2.8$. This is perhaps unsurprising, as CCBS-IBR’s unsoundness is due to how it handles move-wait conflicts; with fewer path options due to a lower average degree, the likelihood that the available paths include a vertex with a stationary agent increases. Over these $11$ problems, CCBS-IBR returns mean $0.243$ and maximum $0.569$ higher sum-of-costs than CCBS-$\delta$-BR. Finally, we note that it is possible for CCBS-IBR— being unsound – to remove all solutions to a problem and therefore never terminate. Such cases are not reflected in our results as we only compare cases where both CCBS-IBR and CCBS-$\delta$-BR return a solution.
We compare the runtime of CCBS-IBR and CCBS-$\delta$-BR in Figure 11 by measuring the number of problems each method is able to solve within a given time limit. To account for different benchmark set sizes, values are normalized to the number of problems CCBS-IBR solves within $30$ seconds. Across all benchmark sets, CCBS-IBR is able to solve more problems within the given time limit. This advantage stems from CCBS-IBR’s more aggressive handling of move-wait conflicts. While this strategy sometimes removes solutions from the search, it often retains at least one optimal solution while reducing the size of the search space. Consequently, CCBS-IBR generally requires fewer node expansions than CCBS-$\delta$-BR, as illustrated in Figure 12. The difference is most pronounced on gridlike roadmaps, followed by roadmaps, suggesting that move-wait conflicts occur more frequently in these environments compared to gridmaps. However, CCBS-IBR’s aggressive pruning comes with trade-offs. In certain cases, as shown above, all optimal solutions are eliminated, leading to sub-optimal results. This means that we are effectively comparing an optimal method (CCBS-$\delta$-BR) to a sub-optimal method (CCBS-IBR). Furthermore, problems where all solutions are eliminated by CCBS-IBR are not reflected in these results, since CCBS-IBR fails to terminate in such cases and contributes to neither the runtime comparison nor the solution quality evaluation above.
Figure 11: The number of solved problems within a given runtime, for each method on each benchmark set, normalized by the method with most solved problems.
Figure 12: Probability density histogram of the increase in the number of high-level node expansions using CCBS-$\delta$-BR compared to CCBS-IBR. The y-axes are limited to maximum $0.8$; the height of the $(1.0,1.1)$-bins are shown specifically.
6.3 Ablation Study on Parameter $\gamma$
The effect of the parameter $\gamma$ — used to compute $\delta$ in CCBS-$\delta$-BR (see Definition 5.1) — on solver runtime is investigated by solving a set of benchmark problems with $\gamma\in\left\{0.1,0.3,...,0.9\right\}$ and recording the number of problems solved within a given time limit. The benchmark set contains gridlike roadmaps: for every average degree value $2.0,2.1,\dots,3.5$, we generate $5$ maps, each with $5$ scenarios. To avoid selection bias, this benchmark is separate from the gridlike roadmap set used in Section 6.2. The results are reported in Figure 13, showing the number of problems that can be solved within $30$ seconds for each value of $\gamma$ and average degree value. We note two things: first, larger $\gamma$ values lead to more problems being solved. Therefore, on these gridlike roadmaps, larger values of $\gamma$ reduce the solver runtime. Second, we see that more problems can be solved within the given time limit on maps with a larger average vertex degree. This is likely due to the larger number of possible paths that an agent can take from its start to goal vertex, making these problems less constrained. At these higher average vertex degree values, the different between the performance using different $\gamma$ values becomes more prominent when compared to lower vertex degrees.
Figure 13: The number of problems solved within a $30$ second time limit for various values of $\gamma$, shown over a range of average vertex degrees on gridlike roadmaps.
7 Conclusions
@@
This work set out to resolve the recently revealed long-standing gap between the theory and practice of Continuous-time Conflict Based Search (CCBS) for optimal Multi-Agent Path Finding in continuous time (MAPFR). We revisited the MAPFR formulation, introduced a new analytical framework for CCBS-style algorithms, and analyzed the publicly available reference implementation of CCBS. The framework provided simple, sufficient conditions for soundness and solution completeness, revealing that the reference implementation violates these conditions — thereby supporting recent counterexamples that challenge the validity of CCBS.
To address this, we proposed a new branching rule ($\delta$-BR) and proved that CCBS using $\delta$-BR satisfies all conditions: it preserves all solutions and guarantees termination within a finite number of iterations on any solvable MAPFR problem. Thus, CCBS using $\delta$-BR is both sound and solution complete for sum-of-costs, makespan, and any other objective function that is strictly monotonically increasing with respect to the maximum agent arrival time. To our knowledge, this represents the first method for MAPFR with the same guarantees as Conflict Based Search for the discrete-time MAPF problem. We illustrated the difference between standard CCBS and CCBS using $\delta$-BR with a counterexample, showing that standard CCBS returns a worse solution than CCBS with $\delta$-BR, thereby showing that standard CCBS is not sound. Furthermore, we compared the two approaches on benchmark problems. While standard CCBS often finds solutions faster due to its more aggressive pruning of the search space, this comes at the cost of discarding all optimal solutions in some cases. In contrast, our branching rule guarantees solution optimality without sacrificing termination.
Finally, we show with our provided implementation that $\delta$-BR can be adopted as a drop-in replacement in existing CCBS implementations. Moreover, the analytical framework and non-termination criterion introduced in this work provide analytical tools for reasoning about CCBS-like solvers and their extensions, thereby providing a rigorous foundation for the next generation of continuous-time multi-agent path finding algorithms.
8 CRediT Authorship Contribution Statement
Alvin Combrink: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing - Original Draft. Sabino Francesco Roselli: Software, Writing - Review & Editing, Supervision. Martin Fabian: Writing - Review & Editing, Supervision, Funding acquisition.
9 Declaration of Competing Interest
Alvin Combrink reports financial support was provided by Sweden’s Innovation Agency. Sabino Francesco Roselli reports financial support was provided by Sweden’s Innovation Agency. Martin Fabian reports financial support was provided by Sweden’s Innovation Agency. If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
10 Acknowledgments
We gratefully acknowledge the Vinnova project CLOUDS (Intelligent algorithms to support Circular soLutions fOr sUstainable proDuction Systems), and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Implementation Details
Our implementation of CCBS-IBR  [^4] https://github.com/Adcombrink/Optimal-Continuous-CBS/tree/originalCCBS closely follows that of the publicly available CCBS implementation [^5] https://github.com/PathPlanning/Continuous-CBS from [21]. However, minor modifications have been made to ensure a fair comparison between CCBS-IBR and CCBS-$\delta$-BR  [^6] https://github.com/Adcombrink/Optimal-Continuous-CBS. For more details than what we describe here, we refer the reader to the provided repositories.
First, we do not use any high-level heuristics or disjoint splitting for CCBS-IBR or CCBS-$\delta$-BR, as these are not the focus of this work. Therefore, ensuring compatibility of CCBS-$\delta$-BR with these additional improvements are left for future work. We do however enable Cardinal Constraints.
Second, in the implementation from [21], move-wait conflict detection is done using the same method as for move-move conflicts: a binary interval search with a specified precision. CCBS-$\delta$-BR, however, is highly sensitive to small numerical differences between the detection of move-wait conflicts and the computation of $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. Specifically, for a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ for which $\delta=t^{j}+w_{D}^{j}-\bar{t_{1}^{c}}$ (see Definition 5.1), the resulting constraint interval using in the constraints on $j$ is $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w_{D}^{j},\bar{t_{2}^{c}})$. This interval overlaps with $j$’s wait action time interval $t^{j},t^{j}+w_{D}^{j}]$ by only a singular time instant, $t^{j}+w_{D}^{j}$. Thus, if the constraint interval is subject to numerical imprecision such that it is computed to $t^{j}+w_{D}^{j}+\epsilon,\bar{t_{2}^{c}})$ for some small $\epsilon>0$, then the resulting constraints using this interval will not forbid $\langle w^{j},t^{j}\rangle$ and therefore the detected move-wait conflict may occur again. Our solution to this is straight-forward: we use $\bar{I^{c}}$ for both constraint creation and move-wait collision detection. An alternative could be to use a fixed precision value as additional margin on the constraints, however, this could be regarded as a form of discretization which we avoid. Besides, since $\bar{I^{c}}$ is computed analytically (under the assumption that edges are traversed in straight lines, as in all experiments), the collision detection is not subject to a precision value and is therefore more precise. Moreover, for the case of general motion functions, CCBS-$\delta$-BR’s numerical sensitivity is easily overcome by using the same computations for collision detection as for constraint creation (such as by pre-computing intersection intervals with, e.g., a binary interval search). Therefore, these modifications do not require agents to only traverse edges in straight lines. As a consequence of using $\bar{I^{c}}$ in the collision detection for CCBS-$\delta$-BR, the same collision detection is used in CCBS-IBR to avoid any differences in the underlying problem that these two methods solve.
Finally, the computation of $\bar{I^{c}}$ is re-implemented for more robustness to numerical imprecision, particularly for near-tangent collisions where the collision interval is singular or close to singular.
Appendix B Validating the Optimality of CCBS-$\delta$-BR’s Solution
We validated the optimality of CCBS-$\delta$-BR’s solution to the counterexample in Section 6.1 using a tailor-made Satisfiability Modulo Theory (SMT) model with the SMT solver Z3 [33]. After roughly $26$ hours of computation, the model returned an equivalent solution with sum-of-costs $9.0$, thereby confirming that CCBS-$\delta$-BR’s solution is an optimal solution. The exact implementation and results can be found in our repository [^7] https://github.com/Adcombrink/Optimal-Continuous-CBS.
The model requires specifying an exact number of timed actions that each agent executes. Thus, the question arises: how many timed actions must be set to ensure that the returned solution is optimal with respect to the MAPFR problem? In the following, we will establish that any solution with a given sum-of-costs requires at most a certain number of timed move actions; if any more timed move actions are required by a single agent, then the sum of costs of such a solution is necessarily higher than the given sum-of-costs. We find for this specific problem that any solution requiring more than $5$ timed move actions necessarily has a sum-of-costs >9.0>9.0. Considering that agents may also execute timed wait actions, we find that the lower bound on the number of timed actions is $11$. Furthermore, we motivate that there does not exists an upper bound to the number of timed actions. Thus, with these results, it holds that the retained solution (which is equivalent to the solution from CCBS-$\delta$-BR) is optimal.
Let $d^{\mathit{LB}}_{i}$ denote an optimistic lower bound on the duration of a plan taking agent $i$ from $S(i)$ to $G(i)$, which is obtained by not considering potential collisions with other agents:
$d^{\mathit{LB}}_{1}=2.5,\qquad d^{\mathit{LB}}_{2}=0,\qquad d^{\mathit{LB}}_{3}=2.0,\qquad d^{\mathit{LB}}_{4}=1.0.$
For solution $\Pi$, let $n$ be the maximum number of timed move actions contained in any single plan $\pi\in\Pi$. Since the minimum move action duration over all move actions in the counterexample is $1.0$, a plan with $n$ move actions has duration $\geq n$. The sum-of-costs lower bound $\mathit{SOC}^{\mathit{LB}}_{\Pi}$ for $\Pi$ occurs when one agent executes $n$ moves while others follow their optimistically shortest-duration plans:
$\mathit{SOC}^{\mathit{LB}}_{\Pi}$ $=\min_{i=1,..4}\left\{n+\sum_{j\in\{1,..,4\}\setminus\left\{i\right\}}d^{\mathit{LB}}_{j}\right\}=\min\left\{\begin{gathered}n+d^{\mathit{LB}}_{2}+d^{\mathit{LB}}_{3}+d^{\mathit{LB}}_{4}=n+3.0\\ n+d^{\mathit{LB}}_{1}+d^{\mathit{LB}}_{3}+d^{\mathit{LB}}_{4}=n+5.5\\ n+d^{\mathit{LB}}_{1}+d^{\mathit{LB}}_{2}+d^{\mathit{LB}}_{4}=n+3.5\\ n+d^{\mathit{LB}}_{1}+d^{\mathit{LB}}_{2}+d^{\mathit{LB}}_{3}=n+4.5\end{gathered}\right\}=n+3.0.$
From this, if any solution has sum-of-costs <9.0<9.0, then $n+3.0<9.0$, implying $n\leq 5$. Since consecutive wait actions can be merged, an agent executing $n$ move actions needs at most $n+1$ wait actions. Therefore, the maximum number of timed actions per agent is $n+n+1=2n+1=11$. Moreover, since wait actions can be split arbitrarily, enforcing that each agent executes exactly $11$ timed actions still allows for an optimal solution to be found.